OpenStudy (anonymous):

A local hamburger shop sold a combined total of 711 hamburgers and cheeseburgers on Wednesday. There were 61 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Wednesday?

OpenStudy (anonymous):

@aum

OpenStudy (justin_lee):

Take the letter H to represent the number of hamburgers sold. Take the letter C to represent the number of cheeseburgers sold. You know that the combined total amount of burgers (both hamburgers and cheesebugers) sold is 711. This means that: H + C = 711 burgers. You also know that there are 61 more cheesebugers sold than hamburgers. Therefore, the amount of cheeseburgers there are are the number of hamburgers plus 61: C = H + 61 extra cheeseburgers. Can you figure out what to do from here? If you need more help, let me know.

OpenStudy (anonymous):

@Justin_Lee IS IT 772

OpenStudy (justin_lee):

It's not 772. Did you get that by adding 711 to 61?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Justin_Lee

OpenStudy (justin_lee):

Remember that 711 is the total amount of burgers, so how can you have 772 cheeseburgers, when the total amount of burgers (both cheesebugers and hamburgers added together) is equal to 711?

OpenStudy (anonymous):

Ok. So what is the next step

OpenStudy (anonymous):

@Justin_Lee

OpenStudy (justin_lee):

Remember: Equation #1: H + C = 711 Equation #2: C = H + 61 Plug equation #2 into equation #1. Solve for H. Then use the value of H you found to solve for C.